Three masses are located in the vertices of an equilateral triangle. Calculate the magnitude and direction of the gravitational force on the mass m₁. Given: m₁ = 38 kg, m₂ = 340 kg, m₃ = 340 kg, r = 38 m. G = 6.674×10–11 N·m²/kg². 


1
Expert's answer
2021-03-21T11:26:03-0400

Let "r" is the side of the triangle. So, we have


"F_2=F_3=G\frac{m_1m_2}{r^2}=6.674\cdot10^{-11}\cdot \frac{38\cdot 340}{38^2}\approx5.97\cdot10^{-10}\ (N)"


"F_{total}=2\cdot5.97\cdot10^{-10}\cdot\cos30°=1.03\cdot10^{-9}\ (N)" . Answer